OpenStudy (anonymous):

Kersha has two jobs. During the day, she works as an office clerk, and in the evening she works as a cashier. Her office job pays her $12.00 per hour, and her cashier job pays her $8.25 per hour. In one week, Kersha worked 55 hours, and she earned a total of $585. Find the number of hours Kersha worked in each job. (Points : 1) Office clerk: 20 hours; cashier: 35 hours Office clerk: 35 hours; cashier: 20 hours Office clerk: 28 hours; cashier: 27 hours Office clerk: 31 hours; cashier: 24 hours

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

Hello :) why don't we try to make equations from the information given to us? (office hours) + (cashier hours) = 55 hours and (amount per hr)*(office hour) + (amount per hr)*(cashier hr) = 585 if you would like to use variables o + c = 55 12o + 8.25c = 585 do you know how to use system of equations to solve from there? :)

OpenStudy (anonymous):

A?

jigglypuff314 (jigglypuff314):

mmm right numbers but not matched with the right job I think

OpenStudy (anonymous):

b?

jigglypuff314 (jigglypuff314):

yup, that's what I got

OpenStudy (anonymous):

Thanks

jigglypuff314 (jigglypuff314):

Glad I could help :)